Need to simplify: the square root of (x-3)squared + (y-5)squared = square root of (x-7)squared + (y+1)squared
Answer by edjones(7311)

x-3+y^2-10y+25=x-7+y^2+2y+1
x-3+y^2-10y+25-x+7-y^2-2y-1=0
-3-10y+25+7-2y-1=0
-12y+28=0
-12y=-28
y=7/3
Ed